A Complete Bibliography of Publications in Limnology and Oceanography: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

14 [Ano59c]. ^{32} [OKB58]. ^{45} [BTH57]. ^{90} [BTE58]. _{12} [Ben57, BB56b, SJM57, SMF57].

-Active [SJM57].

1100-Meter [Die59]. 1952 [Fre58a]. 1956 [Wri58].

Abundance [ER57]. Accompanying [Zob59]. According [Ano58d]. Accumulation [BTE58, Ric56, RW59]. Acids [Sha57, SBM59]. Activation [FM59]. Active [JJ59b, SJM57]. Activity [ER57, Hen59, Mcn59, SMF57]. actuel [Woo58]. Adaptation [RM59]. Added [Hep58]. Adjustment [Rei57]. adults [Ric58b]. aequalis [BB57]. Affecting [JH56, PC56, YD56]. africains [Woo58]. African [Woo58]. Age [MOP59]. Agency [Ano59c]. Agents [JJ59b]. Aggregates [JJ59b]. Air [Mcc59]. Albacore [Mur59]. Algae [Mcl57, OKB58, PT58, Ric56, RW59]. Algal [Pom59, Raw56, Pro57].

Allman [Ano57b]. Along [BB56b]. American [Ano57a, Ket58, Ano57f, Ano59a, Ano59b, Ket59, Ray59]. americana [Hul57]. Amino [SBM59]. Amphipleura [Lew58]. Analyses [NCB59]. Analysis [FM59]. Animal [San58]. Animals [CH59]. Announcement [Red56]. Annual [Man58]. Antarctic [Mcn59]. Antibacterial [Mcn59]. antibiosis [Pro57]. Applications [RC56]. appraisal [Fre57b]. Approach [RD58]. Aquatic [Fre57a, Pen56, Fre57c]. Aqueous [Zob59]. Arctic [LBL58, PS58]. Area [St.58]. arenaria [Aye56b]. artedi [JH56]. Artemia [BTE58]. Articles [Ano57a]. Artificial [Bam58, BA57, EAP56]. Ash [MSH58]. ashlandi [CA59]. Aspects [Col58]. Association [Ano59g].

B [Ben57, BB56b, SJM57, SMF57]. Bacteria [HP58, HA59, JJ59b, RMI58, SJ57, SJM57]. Bacterial [Eme58, JJ59a]. Balance [FNCR56]. balanoides [BB58a]. Balanus [BB58a]. Bank [PN58]. Bar [CB59]. Barnstable [Aye59b]. Basin [Cur57]. Basins [Eme56]. Bathymetrical [Gor58b]. Bathypelagic [HE56]. Bathyscaph [Die59]. Bay [BS59, Wie59b, Hic59, Pra59, San58, Sma57]. Bd [Fre58a]. Beaches [Wie59a]. Behavior [Chr57]. Bell [Ano59h]. Benthic [HB56, OKB58, San58]. Bermuda [Bar59]. Between [MSH58, Mcc59]. Bicarbonate [Man58]. Bight [SB58]. Bioactivity [Ohl56]. Bioassay [Eys58]. Biochemical [Col58]. Biologic [Rei57]. Biological [Ano59g, BA57, Sha57]. Biologists [Ano59a]. Birge [Fre57b]. Bloom [RP57]. Blooms [PTR56]. Bluegill [WJ59]. Boden [BB57]. Body [RT56, SA56]. Bog [SBM59, ZBH56]. Book [Fre58a, Mor59]. Bottle [PB59]. Bottom [And59, Eme58]. Breakdown [Gor58a]. Britain [Gor58b]. British [Fre58b, Ric58b]. Brown [EG57]. Budget [SA56, Gei57]. Bumpus [YD56, CC57]. Buzzards [San58]. By-Laws [Ano57a].

C [Ano59c]. caddisflies [Ric58a]. Cairn [Gor57b]. Calanoid [CC57]. Calcium [BTH57, Car57]. Calibration [YD56]. California [Eme56, Fre57a]. Canadian [HA58, PS58]. Canyon [Wri58, Wri59]. Cape [St.58]. Carbohydrates [GW58]. Carbon [BO59]. Carolina [Pie58]. Carotenoids [Val56]. carpio [WP57]. Catch [MOP59]. Catches [Mur59]. caudata [Ano57g]. Cayuga [Hen59]. Cedar [SBM59]. Central [Ano59c, Ril57]. Cerium [RW59]. Cerium-144 [RW59]. Certain [Mcl57]. Chaetognatha [Bie59, Pie58]. Chagrin [Ano57b]. Chain [Sai56]. Chamber [Lun59]. Change [Far59, FLM58]. Changes [WJ59, Zob59]. Characteristics [And58a, Car58, HA58]. Charlottenlund [Ano59c]. Chemical [Car58, Gor57a, Gor57b, Sha57]. Chemistry [HA58, Mor59]. Chesapeake [Wie59b]. Cheshire [Gor57c]. Chlamydomonas [Pro57]. Chlorophyll [MS59, Mcl57, RY57, Shi58, YR57]. Chlorophyll-a [Mcl57]. Choice [Wie56]. Cinematography [Bay59]. cinerea [Ano57g]. Circularly [Mcl57]. Circulation [CB59]. Cisco [JH56]. Cladocera [Fre58b]. Clam [Aye56b]. Clarity [Mur59]. Clarke [CC57, YD56]. Clarke-Bumpus [YD56].

Classification [Ano58d, HA59, Ric58a]. Coast [BB56b, Mil58, Rei57, HA58]. Cobalt [Ben57]. Coelenterates [HE56]. Cold [KC59]. Coliform [RMI58]. Collected [BB56b]. Color [FLM58]. Comment [Wha58]. Communication [Ano59c]. Communities [Pen57]. Community [Odu57]. Comparison [FM59, Rei58]. Complete [CRB59]. Composition [Gor57a, Gor57b, Gor57c, MOP59, Pen57]. Compression [Zob59]. Computations [HE57]. Concentrates [Kut58]. Concentration [MM59]. Concentrations [Bar57]. concerning [Ano59c]. Concretion [Eys58]. Concretion-Forming [Eys58]. Condition [EG57]. Conditions [CRB59, HP58, KC59, Zob59]. Conference [Ano59d]. Congress [Ano58c]. connaissances [Woo58]. Construct [BO59]. Consumed [HM59]. Content [Shi58, Tur57]. Continental [RY58]. Control [You56]. Controlling [WP57]. Conventional [FM59]. Conversion [Ano57c, MM59]. Copepod [CC57]. Copepods [Con59, Ray59]. Copper [SJ57]. Coral [KH57]. Cordylophora [Ano57b]. Core [Mac58]. Cores [HM59, Tur57]. Correction [Dav59a]. Counting [Lun59]. County [Gor57a]. Crabs [FLM58]. Creek [SBM59]. Cromwell [Ano59h, Mon59]. Crop [Bar59, HA59, Wri59]. Crustacea [Wie56]. Crustaceans [BB57]. Cryptochironomus [Cur58]. Cullinectes [CRB59]. Culture [MM59, RA58]. Cumacea [Wie56]. Cumella [Wie56]. Current [HE57, Woo58]. Currents [Aye56a]. Curves [BO59]. Cycles [CA59]. Cyprinus [WP57].

Daily [DO57]. Daphnia [SR56]. Dark [PB59]. Data [RY57]. Daytime [Ver57]. Deep [Aye56a, Eme56, Rei58, Tur57, Zob59]. Deep-Sea [Zob59]. Delaware [Hul57]. Denmark [Ano59c]. Deposits [Mac58]. Depths [CH59]. Description [CC57]. Determination [Aye56a, Car57, HE57, MM59, Sut57, Ano59c]. Determinations [RC56, Ste58]. Determined [JJ59a, JJ59b]. Deuterium [FNCR56]. deutschen [Fre58a]. Development [BB58a, CA59, CRB59, LBL58]. Device [Ano57c]. Diaptomids [RT56]. Diaptomus [CA59]. Different [JJ59a]. Dinoflagellate [PTR56, RP57]. Dioxide [BO59]. Diptera [Cur58]. Dismal [SBM59]. Disposal [Ano59d]. Distribution [Bie59, BB57, CC57, Cur59, Hul57, St.58, SBM59, Wie59a, Wil58]. District [Gor57b]. Disturbed [Rei57]. Diurnal [Shi58]. Dive [Die59]. Donax [TB57]. Dr. [Ano59h]. Drogue [Rei58]. drop [BB56a]. Due [Col59]. during [KC59, Wri58]. Dynamic [Aye56a, HE57]. Dynamics [Aye56b, Hep58, Rou59a, Rou59b].

E. [Fre57b]. Early [Far59]. East [HA58, Woo58]. Eastern [Shi58]. Ecological [Bam58]. Effect [CG59, MSH58, Mcl57, Mur59, SR56, SJ57, Wie59a]. Effects [ER57, LBL58, Mil58]. Egg [Dav59b]. Eggs [JH56, RT56]. Electrical [WP57]. Electrode [Kan59]. Eleven [Odu57]. Energiehaushalt [Gei57]. Energy [Ohl56, WP57, Gei57]. England [Gor57c, Mil58]. Enumeration [JJ59a]. Environment [LBL58]. Environmental [FLM58, PC56]. Environments [SJ57]. Epiphasic [Val56]. Epitome [Gor58b]. Equilibrium [BTE58, HP58]. Ergebnisse [Fre58a]. Erie [Cur57, Cur59]. Errata [Ano58a, Mon58]. Erratum [Ano57d]. Error [PB59]. Esox [Sai56]. Esquire [Ano59h]. est [Woo58]. est-africains [Woo58]. Estimates [Kut58, Sai56]. Estimation [RY57]. Estuaries [AI58]. Estuary [Hul57, Rei57]. état [Woo58]. Euphausiid [BB57]. Euphotic [Wri58]. Eupleura [Ano57g]. Eutrophication [EAP56]. Evasterias [Chr57]. Evidence [DO57, Hen59]. Evolution [Ric58a]. Exchange [Ric56]. Expedition [Fre58a]. Exploited [Rou59a, Rou59b]. explorer [Fre57b]. Extracellular [GW58]. Extracts [WV57].

Factor [Wil58]. Factors [BA57, CG59, FLM58, JH56, MM59, PC56, WP57, Wie56, YD56, You56]. Fauna [And59]. Features [And58b]. Feeding [Chr57]. Fellowships [Ano59e]. Ferry [Wri58, Wri59]. Fertilization [ER57]. Fertilized [Bar57]. Fiddler [FLM58]. Field [Col59, WP57]. Fish [Bar59, BTH57, Col59, CG59, Dav59b, HB56, Rou59a, Rou59b, WB57, WP57, BB58b]. Fishery [Ano59a]. Fishes [Far59, HHWB58, Sul58]. Fishponds [Hep58]. Fixed [SR56]. Flagellates [GW58]. Flashing [CH59]. Florida [Odu57, You56]. Flotation [And59]. Flowing [Odu56]. Fluctuation [Shi58]. Fluctuations [Man58]. Food [EG57]. Forbes [SR56]. Formation [Gor58a]. Forming [Eys58]. Foundation [Ano59i]. Four [Far59]. Free [WV57]. French [Woo58]. Fresh [Ben57, Dav59b]. Freshwater [Gor58b, LBL58, WB57, Fre58b]. Function [Ryt56b].

GEK [Rei58]. Geography [Mor59]. Geology [Aye59a]. Georges [PN58]. Georgia [BB56b, Pom59]. German [Gei57]. Glacial [Val56]. Gmelin [PN58]. Gorm [Gor57b]. Grain [Wie59a]. grands [Woo58]. Grants [Ano59e, Ano59k]. Grass [Odu57]. gravity [BB58b]. Great [CH59, Woo58, Aye59a]. Growth [Far59, PC56, PT58, SJ57]. Gulf [Rei57, SB58].

Haematococcus [Pro57]. Halifax [Gor57a]. Handbuch [Ess57]. Hansen [BB57]. Harbor [Ano57b, Aye59b]. Hardness [Man58]. Hart [Wie56]. Hatching [JH56]. Hatteras [St.58]. Hawaiian [KH57]. Heat [SA56]. Height [Aye56a, HE57]. History [RP57]. Homing [HHWB58]. Hot [Pat58]. Huron [BS59]. Hydrobotanik. [Gei57]. Hydrobotany [Gei57]. Hydrographic [Rei57]. Hydrography [Aye59b].

II [HRC58, Wri59]. III [BTH57, MSH58, Ray59]. Improved [Sut57]. Inc. [Ano57a, Ano57f, Ket58, Ket59]. Incorporation [Ano57a]. Index [Ano57e, Ano58b, Ano59f]. India [Ano59g]. Indicators [Raw56]. Influence [GP58]. Influenced [HB56]. Influencing [Wie56]. Infra [Bay59]. Infra-red [Bay59]. Inhabiting [Wie59a]. Inorganic [MSH58]. Input [WP57]. insects [Fre57a]. Inshore [Pie58]. Institute [Ano59a]. Intensity [Ryt56b]. Interface [HRC58]. Internal [CC57, Eme56, Hen59]. International [Ano58c]. Invertebrates [Wie59a]. Investigation [CB59]. Involved [FLM58]. Ionic [Gor57c]. Iron [SB58]. Island [BS59, PS58, Sai56]. Isolated [SJ57]. Isopropyl [WB57]. Israel [Hep58]. IV [HP58, KC59].

Jet [BB56a].

Kentucky [KC59]. key [Fre58b, Ric58b]. Knobs [KC59]. knowledge [Woo58].

L. [BB58a, BTE58, WP57]. Laboratory [CRB59]. lacs [Woo58]. lacustris [Ano57b]. Lake [And58b, CB59, Cur59, ER57, Eys58, Gor57a, Gor57c, HA58, HRC58, HP58, HM59, HA59, Hen59, Mac58, MSH58, Raw56, Sha57, Val56, WV57, ZBH56, BB56a, BS59, CA59, Cur57, EAP56, FNCR56, JH56, KC59, Tal57]. Lakes [And58a, Aye56a, Bar57, Gor58a, HE57, HA58, HA59, LBL58, Mcc59, Ohl56, RT56, Ver56, Fre57b, Woo58, Aye59a]. Larvae [BB58a, Col59, Cur58]. Larval [CRB59, Far59]. Laws [Ano57a]. Length [MOP59]. LeSueur [JH56, Sai56]. Leucichthys [JH56]. leur [Woo58]. Levels [Mil58]. Life [RP57]. Light [Mcl57, PB59, Ryt56b, RY57, RM59]. Limit [Sai56]. Limnetic [Pen57]. Limnological [And58b, KC59]. limnologischen [Fre58a]. limnologist [Fre57b]. Limnology [Ano57a, Ano57f, Ano59b, Gor58b, Ket58, Ket59, Wri58, Wri59, Mor59, Ano57d, Ano57e, Ano58a, Ano58b, Ano59f, Mon58]. Linnaeus [EG57]. live [Ano57g]. live-sexing [Ano57g]. Lochs [Gor58b]. Loss [RW59]. Lower [Sma57]. Lowland [Gor57c]. Luminescent [CH59].

M [Ano59h]. magellanicus [PN58]. Manitoulin [BS59]. manual [Fre57c]. Maracaibo [FNCR56]. Marine [Ano58d, Aye56b, BTH57, Col59, Con59, Far59, FM59, GW58, Mcl57, Mcn59, NCB59, Odu57, OKB58, Ric56, RW59, RMI58, RM59, SJ57, SJM57, Ano59g]. Marl [Eys58]. Marsh [BB56b, CA59]. Marshes [Pom59]. Mass [RA58]. Massachusetts [Aye59b]. Masses [Bie59]. Maximum [Sai56]. Md. [Wie59b]. Measurement [BO59, Ryt56a]. Measurements [Ano59c, Odu57, PS58, Rei58]. Mechanism [BS59]. Medial [MM59]. Meer [Wie58]. Meere [Ess57]. memoir [Fre57b]. Memorial [Ano59h]. Mendota [JH56, BB56a, CB59]. Menhaden [MOP59]. mercenaria [PC56]. Meromictic [And58b]. Metabolism [BTH57, WJ59]. Meter [Die59]. Method [Aye56a, HE57, PB59, Sut57]. Methods [FM59, JJ59a, Ano57g]. Methylphosphonofluoridate [WB57]. Michigan [Cur58, EG57]. Microcrustacea [Bay59]. Microzone [Gor58a]. Midwater [Aro59]. Migrations [TB57]. minax [FLM58]. Minimum [Sai56]. Minnesota [SBM59]. Mixing [BS59]. Mixture [BTE58]. Modified [And59, CC57]. Monochromator [Tyl59]. Mortality [Col59]. Mountain [MS59, RT56, Ric58a]. Mucilage [Lew58]. Mud [Gor58a, HRC58, HP58, MSH58]. Muds [BB56b, ZBH56]. Mullet [SMF57]. Mya [Aye56b].

Nannoplankton [Lun59]. Narragansett [Sma57, Hic59, Pra59]. National [Ano59i]. Natural [Car57, Gor57b, Tyl59]. Nematodes [Wie59b]. Neomysis [Hul57]. Newborn [SR56]. niger [Sai56]. Nile [PT58]. Nitrogen [YV58]. North [Aro59, Pie58, Ril57]. Northern [Gor58b]. nos [Woo58]. Note [CRB59, Wie59b]. Notes [Kut58, Wha58, Ano57g]. Notice [Ano59j]. Nova [Gor57a]. Number [RT56, YD56]. Numbers [You56]. Numerical [Kut58]. Nutrients [MM59]. nymphs [Ric58b].

Observation [Bay59, Col59, PN58]. Observations [Ben57, Gor58a, JH56, PTR56, Tal57, Mon57]. Occurrence [WV57]. Ocean [Ryt56b, RY57]. Oceanic [CH59, Mon57]. Oceanographic [Ano58c, Col58, RC56]. Oceanography [Ano57a, Ano57f, Hic59, Ket58, Ket59, Ano59b, Ano57d, Ano57e, Ano58a, Ano58b, Ano59f, Mon58]. Oceans [YV58]. October [Wri58]. off [RY58]. Offshore [SB58]. Ohio [Ano57b]. Organic [KH57, Sha57]. Organisms [BTH57, FM59]. Orientation [BB58b, HHWB58]. Origin [Cur59, LBL58]. our [Woo58]. Outfalls [RMI58]. Oxidation [HRC58]. Oxidation-Reduction [HRC58]. Oxidized [Gor58a, HP58, MSH58]. Oxygen [GP58, Kan59, PB59, Ray59, RC56, Ste58, HM59]. Ozeane [Ess57].

P [OKB58]. Pacific [Shi58, Aro59, Bie59, BB57, Mon57]. Panama [SB58]. Particulate [SB58]. Partition [MSH58]. Patterns [CC57]. Peats [SBM59]. Pentachlorophenate [CG59]. Percentages [SR56]. Periodicity [DO57]. pH [BO59, MSH58]. Phaeodactylum [RA58]. Phases [FLM58]. Phosphate [MSH58]. Phosphorus [Cur57, Cur59, Hep58, ZBH56]. Photo [DO57]. Photosynthesis [Mcl57, Ryt56b, Ver57]. Photosynthetic [Shi58]. Physical [Gor58b, Hic59, Car58]. physics [Mor59]. physiology [Sul58]. Phytoplancton [Woo58]. Phytoplankton [CA59, GP58, Mcn59, Pra59, Ril57, RY57, RM59, Shi58, Sma57, Ver57, YR57, YV58, Woo58, Woo58, Wri58, Wri59]. Pickerel [Sai56]. Place [HA59]. Placopecten [PN58]. Plankton [CC57, Kut58, NCB59, PT58, Sut57, YD56]. Planktonic [Bam58, Bie59, Dav59b, Ray59, Ric56, RW59]. Plants [HP58, Pen56, Fre57c]. Plecoptera [Ric58b]. Point [CB59]. Polarized [Mcl57]. Polarographic [Kan59]. Pond [Bam58]. Ponds [BA57, Man58]. Population [Aye56b, CC57, CA59]. Populations [JJ59a, Rou59a, Rou59b, SR56]. Portable [Mac58]. Post [Val56]. Post-Glacial [Val56]. Potassium [Bar57]. Precision [Kut58]. Predation* [HB56]. Preface [Fre56]. Preliminary [Ben57, CRB59]. Preserved [Sut57]. previous [Ano57g]. Primary [Fre56, KH57, Odu56, Odu57, OKB58, Pen56, PS58, Ryt56a, RY58, Ver56, Wri59, Ano59c]. Problems [Col58, RD58]. Proceedings [Ano57f, Ket59, Ket58]. Production [Fre56, GW58, Odu56, Odu57, Ohl56, Pen56, PS58, RD58, Ryt56a, RY57, RY58, Sai56, SJM57, Ver56, Wri59, Ano59c]. Productivity [BO59, HB56, KH57, MS59, OKB58, Pom59]. Protozoa [Bam58]. Puget [Wie59a]. pugilator [FLM58]. pulicaria [SR56]. Pump1 [Wha58]. Pupae [Cur58]. Pure [Ste58]. Pushnet [AI58].

Quantitative [AI58, CH59, YD56].

Radioactive [RW59]. Radionuclides [BTH57]. Radiophosphorus [HP58]. Rapid [Ano57c, Ano57g]. Rate [BB58a, Con59, Shi58]. Rates [Far59]. Rathbun [CRB59]. Raw [BA57]. Records [CH59]. Red [Col58, Bay59]. Reduced [HP58]. Reduction [HRC58]. Reef [Bar59, KH57]. Regional [Con59]. Related [Col58, CA59, RT56]. Relations [HRC58, Mcc59]. Relationship [Bie59, EG57]. Relationships [San58, Wri58]. Release [ZBH56]. Removal [SR56]. Renewal [RT56]. Requirements [Ray59]. Research [Ano59a, Ano59e, Ano59k]. Reservoir [Wri58, Wri59]. Resolution [Ano59h]. Respecting [Ano59h]. Respiration [GP58, Ray59]. Respiratory [Con59]. Response [WB57]. Retween [EG57]. Review [Fre58a, Mor59]. Rhode [Sai56]. Rhythm [FLM58]. River [MS59, Hul57]. rutilans [Lew58].

salina [BTE58]. Saline [And58a, And58b]. Salinity [Ano57c, Ano58d]. Salmo [EG57]. Salmon [ER57]. Salt [Pom59]. Salton [Car58]. sample [Kut58]. Sampler [CC57, Eme58, Mac58, YD56]. Samples [And59]. Sampling [AI58, Wha58]. sappers [CRB59]. Sargasso [Ril57]. Sarin [WB57]. Say [TB57]. Scallop [PN58]. Science [Ano59i]. Scotia [Gor57a]. Scotland [Gor57b]. Scottish [Gor58b]. Sea [JJ59a, JJ59b, Pat58, PN58, RC56, Tur57, Zob59, Ano59c, Car58, Ril57]. Seasonal [And58a, CA59, Con59, PT58, WJ59]. Seastar [Chr57]. Second [CB59]. Secondary [RD58]. Sediment [HA58, HRC58, HP58, HM59, HA59, MSH58, San58, WV57]. Sediments [HA59, RMI58, Val56]. September [Wri58]. Serum [SMF57]. Setting [FLM58]. Sewage [BA57, RMI58]. sexing [Ano57g]. Shallow [AI58, And58b, Bar59, JH56]. Shallow-Water [JH56]. Shark [SMF57]. Shelf [RY58]. Shimada [Ano59h]. Short [YR57]. Short-Term [YR57]. Shrimp [AI58, Wil58]. Significance [Tur57, YR57]. Silver [You56]. Simple [Lun59]. Simplified [HE57]. Size [RT56, Sai56, SR56, Wie59a]. Size-Limit [Sai56]. Slot [Ano59c]. Small [Bam58, Kut58, Wie59a]. Small-sample [Kut58]. Society [Ano57a, Ano57f, Ano59b, Ket58, Ket59]. Sodium [CG59]. Solids [BB56b]. Solubility [RC56, Ste58]. Solution [BTH57]. Solutions [Zob59]. Some [And58b, Col58, Gor57b, Gor57c, GW58, HA58, JH56, Mcc59, Ray59, RD58, RC56, Tal57, WB57, YD56, KC59]. Sorting [And59]. Sound [Wie59a]. Source [Cur57]. Sources [PB59]. South [BS59]. Southampton [PS58]. Spawning [PN58]. Species [Cur58, Pen57, Ray59, You56, Fre58b]. Spectrographic [NCB59]. Spectrum [Pat58]. Spey [Gor57b]. Springs [Odu57, You56]. Sr [BTE58]. Standing [Bar59, Eme56, HA59, Wri59]. state [Woo58]. Stimpson [Chr57]. stoneflies [Ric58b]. Strand [Wie58]. Strath [Gor57b]. Stratification [Tal57]. Stratigraphic [SBM59]. Stream [EG57]. Stress [BB56a]. Stress-drop [BB56a]. Strontium [Ric56, Tur57]. Studies [Aro59, Bam58, KC59, Pat58, Pro57, RA58, San58, Sha57, Sma57]. Study [St.58]. subaequalis [BB57]. Submersible [Wha58]. Substances [SJM57]. Substrates [Wil58]. Substratum [Wie56]. Subterranean [Wie59b]. Succession [PT58]. Sugars [WV57]. Sulphate [Man58]. Summer [Bar59]. Sun [HHWB58]. Sun-Orientation [HHWB58]. Supply [EG57]. Surface [Gor58a, JJ59b, Mcc59]. Survey [Gor58b]. Survival [JH56]. Suspended [BB56b]. Swamp [SBM59]. Symposium [Fre56]. Synthesis [Mcl57]. synthetic [DO57]. System [Ano58d].

Taxonomy [BB57]. Technique [And59, Ano57g]. Temperatures [Mcc59]. Tension [GP58]. Term [YR57]. Their [BB57, YR57, Bie59, Woo58]. Thermal [Zob59]. Three [RMI58]. Thysanopoda [BB57]. Tidal [FLM58, TB57]. Tides [Col58]. Titration [Ano57c]. Townsend [Ano59h, Mon59]. Toxicity [CG59]. Trace [FM59]. Travel [Ano59i]. Trawling [Aro59]. treatise [Mor59]. Treatment [BA57, JJ59b]. Trieste [Die59]. Trophic [Raw56]. troschelii [Chr57]. Trout [Bar57, EG57]. trutta [EG57]. Tubes [Lew58]. Turbulence [Pat58]. Turtle [Odu57]. Turtle-Grass [Odu57]. Two [And58a, PB59]. Types [Raw56].

Uca [FLM58]. Unusually [KC59]. Uptake [BTH57, OKB58, RW59]. Urosalpinx [Ano57g]. Use [BO59, MM59]. using [Pro57]. Utilization [Ohl56].

V [HM59]. Values [Ano57c]. variabilis [TB57]. Variability [SR56]. Variation [Con59]. Variations [Tur57, Ver57, YR57]. Various [CG59]. Varying [SA56]. Vascular [Pen56]. Venezuela [Fre58a]. Venezuela-Expedition [Fre58a]. Venice [Ano58d]. Venus [PC56]. VI [HA59]. Victoria [Tal57]. Virginia [MOP59, SBM59]. Vitamin [Ben57, BB56b, SJM57, SMF57]. Vitamin-B [SMF57]. Vol [Ano57d, Ano58a, Mon58]. Volcanic [ER57]. Volume [Mor59, SA56]. Volumes [Sut57]. Volumetric [Kut58, St.58]. vulgaris [Wie56].

Warming [Col59]. Washington [And58a, CA59, EAP56]. Waste [Ano59d]. Water [Ben57, Bie59, Dav59b, Eys58, HA58, HRC58, HP58, HM59, HA59, JJ59a, JH56, JJ59b, MSH58, Mcc59, Mil58, Mur59, RT56, Rei58, RC56, SA56, Sha57, Ste58, Tyl59, WB57]. Waters [Ano58d, BS59, Car57, Gor57a, Gor57b, Gor57c, MOP59, Odu56, Pie58, RY58, SB58, Shi58]. Wave [Hen59]. Waves [Eme56]. Weight [MOP59]. Wellen [Ess57]. Western [Cur57, RD58, Cur59]. White [PT58]. Wild [EG57]. Winds [Mil58]. Winter [KC59]. Wires [Pat58]. Wisconsin [JH56].

Yellow [Sha57]. Yield [Sai56]. York [Hen59, RY58]. Young [JH56].

Zone [Wri58]. Zooplankton [Pen57, St.58, Wri58].

References

Allen:1958:PQS

[AI58] Donald M. Allen and Anthony Inglis. A pushnet for quantitative sampling of shrimp in shallow estuaries. Limnology and Oceanography, 3(2):239–241, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1958:SCT

[And58a] G. C. Anderson. Seasonal characteristics of two saline lakes in Washington. Limnology and Oceanography, 4(2):51–68, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1958:SLF

[And58b] G. C. Anderson. Some limnological features of a shallow saline meromictic lake. Limnology and Oceanography, 3(3):259–270, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1959:MFT

[And59] Richard O. Anderson. A modified flotation technique for sorting bottom fauna samples. Limnology and Oceanography, 4(2):223–225, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1957:AIL

[Ano57a] Anonymous. Articles of incorporation and by-laws of The American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 2(1):51–54, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1957:CLA

[Ano57b] Anonymous. Cordylophora lacustris Allman from Chagrin Harbor, Ohio. Limnology and Oceanography, 1(4):158–159, April 1957. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1957:DRC

[Ano57c] Anonymous. A device for rapid conversion of titration values to salinity. Limnology and Oceanography, 2(4):371–374, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1957:EVL

[Ano57d] Anonymous. Erratum in vol. 1 of Limnology and Oceanography. Limnology and Oceanography, 2(2):154, April 1957. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1957:ILO

[Ano57e] Anonymous. Index to Limnology and Oceanography. Limnology and Oceanography, 2(4):375–380, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1957:PAS

[Ano57f] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 2(1):47–50, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1957:RLS

[Ano57g] Anonymous. A rapid live-sexing technique for Urosalpinx cinerea and Eupleura caudata, with notes on previous methods. Limnology and Oceanography, 2(1):41–42, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1958:EVL

[Ano58a] Anonymous. Errata in vol. 2 of Limnology and Oceanography. Limnology and Oceanography, 4(2):118, January 1958. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1958:ILO

[Ano58b] Anonymous. Index to Limnology and Oceanography. Limnology and Oceanography, 3(4):484–488, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1958:IOC

[Ano58c] Anonymous. International oceanographic Congress. Limnology and Oceanography, 3(3):347–348, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1958:VSC

[Ano58d] Anonymous. The Venice system for the classification of marine waters according to salinity. Limnology and Oceanography, 3(3): 346–347, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:AIF

[Ano59a] Anonymous. American institute of fishery research biologists. Limnology and Oceanography, 4(4):503, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:ASL

[Ano59b] Anonymous. American Society of Limnology and Oceanography. Limnology and Oceanography, 4(3):357–365, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:CCC

[Ano59c] Anonymous. Communication concerning a central agency for C^{14} determination (measurements of primary production in the sea) at Charlottenlund Slot, Denmark. Limnology and Oceanography, 4(1): 106–107, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:CWD

[Ano59d] Anonymous. Conference on waste disposal. Limnology and Oceanography, 4(2):228–229, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:FRG

[Ano59e] Anonymous. Fellowships and research grants. Limnology and Oceanography, 4(2):228, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:ILO

[Ano59f] Anonymous. Index to Limnology and Oceanography. Limnology and Oceanography, 4(4):504–507, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:MBA

[Ano59g] Anonymous. Marine Biological Association of India. Limnology and Oceanography, 4(4):503, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:MRR

[Ano59h] Anonymous. Memorial resolution respecting Dr. Bell M. Shimada and Townsend Cromwell, Esquire. Limnology and Oceanography, 4 (4):502, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:NSF

[Ano59i] Anonymous. National Science Foundation travel. Limnology and Oceanography, 4(1):106, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:N

[Ano59j] Anonymous. Notice. Limnology and Oceanography, 4(2):229, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1959:RG

[Ano59k] Anonymous. Research grants. Limnology and Oceanography, 4(4): 503, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Aron:1959:MTS

[Aro59] William Aron. Midwater trawling studies in the North Pacific. Limnology and Oceanography, 4(4):409–418, April 1959. CODEN LI-OCAH. ISSN 0024-3590.

Ayers:1956:DHM

[Aye56a] John C. Ayers. A dynamic height method for the determination of currents in deep lakes. Limnology and Oceanography, 1(3):150–161, July 1956. CODEN LIOCAH. ISSN 0024-3590.

Ayers:1956:PDM

[Aye56b] John C. Ayers. Population dynamics of the marine clam, Mya arenaria. Limnology and Oceanography, 1(1):26–34, January 1956. CO-DEN LIOCAH. ISSN 0024-3590.

Ayers:1959:GGL

[Aye59a] John C. Ayers. Geology of the Great Lakes. Limnology and Oceanography, 4(3):356, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Ayers:1959:HBH

[Aye59b] John C. Ayers. The hydrography of Barnstable Harbor, Massachusetts. Limnology and Oceanography, 4(4):448–462, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Bartsch:1957:BFT

[BA57] A. F. Bartsch and M. O. Allum. Biological factors in treatment of raw sewage in artificial ponds. Limnology and Oceanography, 2(2): 77–84, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Bamforth:1958:ESP

[Bam58] Stuart S. Bamforth. Ecological studies on the planktonic protozoa of a small artificial pond. Limnology and Oceanography, 3(4):398– 412, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Barrett:1957:PCF

[Bar57] Paul H. Barrett. Potassium concentrations in fertilized trout lakes. Limnology and Oceanography, 3(1):287–294, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Bardach:1959:SSC

[Bar59] John E. Bardach. The summer standing crop of fish on a shallow Bermuda reef. Limnology and Oceanography, 4(1):77–85, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Baylor:1959:IRO

[Bay59] Edward R. Baylor. Infra-red observation and cinematography of microcrustacea. Limnology and Oceanography, 4(4):498–499, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Bryson:1956:SDJ

[BB56a] Reid A. Bryson and W. W. Bunqe, Jr. The “Stress-drop jet” in Lake Mendota. Limnology and Oceanography, 1(1):42–46, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Burkholder:1956:VBS

[BB56b] Paul R. Burkholder and Lillian M. Burkholder. Vitamin B_{12} in suspended solids and marsh muds collected along the coast of Georgia. Limnology and Oceanography, 1(3):202–208, July 1956. CODEN LIOCAH. ISSN 0024-3590.

Boden:1957:ECT

[BB57] Brian P. Boden and Edward Brinton. The euphausiid crustaceans Thysanopoda aequalis Hansen and Thysanopoda subaequalis Boden, their taxonomy and distribution in the Pacific. Limnology and Oceanography, 2(4):337–341, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1958:RDB

[BB58a] H. Barnes and Margaret Barnes. The rate of development of Balanus balanoides (L.) larvae. Limnology and Oceanography, 4(2): 29–32, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Braemer:1958:OFG

[BB58b] Wolfgang Braemer and Helga Braemer. Orientation of fish to gravity. Limnology and Oceanography, 3(4):362–372, October 1958. CO-DEN LIOCAH. ISSN 0024-3590.

Benoit:1957:POC

[Ben57] Richard J. Benoit. Preliminary observations on cobalt and vitamin B_{12} in fresh water. Limnology and Oceanography, 3(1):233–240, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Bieri:1959:DPC

[Bie59] Robert Bieri. The distribution of the planktonic Chaetognatha in the Pacific and their relationship to the water masses. Limnology and Oceanography, 4(1):1–28, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Beyers:1959:UCD

[BO59] Robert J. Beyers and Howard T. Odum. The use of carbon dioxide to construct pH curves for the measurement of productivity. Limnology and Oceanography, 4(4):499–502, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Bryson:1959:MMW

[BS59] Reid A. Bryson and Charles R. Stearns. A mechanism for the mixing of the waters of Lake Huron and South Bay, Manitoulin Island. Limnology and Oceanography, 4(3):246–251, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Boroughs:1958:AEM

[BTE58] Howard Boroughs, Sidney J. Townsley, and Winifred Ego. The accumulation of Y^{90} from an equilibrium mixture of Sr^{90}–Y^{90} by Artemia salina (L.). Limnology and Oceanography, 3(4):413–417, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Boroughs:1957:MRM

[BTH57] Howard Boroughs, Sidney J. Townsley, and Robert W. Hiatt. The metabolism of radionuclides by marine organisms. III. The uptake of calcium^{45} in solution by marine fish. Limnology and Oceanography, 2(1):28–32, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Comita:1959:SDP

[CA59] G. W. Comita and G. C. Anderson. The seasonal development of a population of Diaptomus ashlandi Marsh, and related phytoplankton cycles in Lake Washington. Limnology and Oceanography, 4(1): 37–52, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1957:DCN

[Car57] James H. Carpenter. The determination of calcium in natural waters. Limnology and Oceanography, 3(1):271–280, July 1957. CO-DEN LIOCAH. ISSN 0024-3590.

Carpelan:1958:SSP

[Car58] Lars H. Carpelan. The Salton Sea. Physical and chemical characteristics. Limnology and Oceanography, 3(4):373–386, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Clarke:1959:ICS

[CB59] David B. Clarke and Reid A. Bryson. An investigation of the circulation over Second Point Bar, Lake Mendota. Limnology and Oceanography, 4(2):140–144, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Comita:1957:IDP

[CC57] Gabriel W. Comita and Jean J. Comita. The internal distribution patterns of a Calanoid copepod population, and a description of a modified Clarke–Bumpus plankton sampler. Limnology and Oceanography, 2(4):321–332, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Crandall:1959:EVF

[CG59] Catherine Ann Crandall and Clarence J. Goodnight. The effect of various factors on the toxicity of sodium pentachlorophenate to fish. Limnology and Oceanography, 4(1):53–56, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Clarke:1959:QRL

[CH59] George L. Clarke and Charles J. Hubbard. Quantitative records of the luminescent flashing of oceanic animals at great depths. Limnology and Oceanography, 4(2):163–180, April 1959. CODEN LI-OCAH. ISSN 0024-3590.

Christensen:1957:FBS

[Chr57] Aage Møller Christensen. The feeding behavior of the seastar Evasterias troschelii Stimpson. Limnology and Oceanography, 2(3): 180–197, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Collier:1958:SBA

[Col58] Albert Collier. Some biochemical aspects of red tides and related oceanographic problems. Limnology and Oceanography, 4(2):33–39, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Colton:1959:FOM

[Col59] John B. Colton, Jr. A field observation of mortality of marine fish larvae due to warming. Limnology and Oceanography, 4(2):219–222, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Conover:1959:RSV

[Con59] Robert J. Conover. Regional and seasonal variation in the respiratory rate of marine copepods. Limnology and Oceanography, 4(3): 259–268, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Costlow:1959:PNC

[CRB59] John D. Costlow, Jr., George H. Rees, and C. G. Bookhout. Preliminary note on the complete larval development of Cullinectes sappers Rathbun under laboratory conditions. Limnology and Oceanography, 4(2):222–223, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Curl:1957:SPW

[Cur57] Herbert Curl, Jr. A source of phosphorus for the western basin of Lake Erie. Limnology and Oceanography, 2(4):315–320, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Curry:1958:LPS

[Cur58] Laverne L. Curry. Larvae and pupae of the species of Cryptochironomus (Diptera) in Michigan. Limnology and Oceanography, 3 (4):427–442, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Curl:1959:ODP

[Cur59] Herbert Curl, Jr. The origin and distribution of phosphorus in Western Lake Erie. Limnology and Oceanography, 4(1):66–76, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Davis:1959:C

[Dav59a] Charles C. Davis. Correction. Limnology and Oceanography, 4(2): 229, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Davis:1959:PFE

[Dav59b] Charles C. Davis. A planktonic fish egg from fresh water. Limnology and Oceanography, 4(3):352–355, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Dietz:1959:MDB

[Die59] Robert S. Dietz. 1100-meter dive in the bathyscaph Trieste. Limnology and Oceanography, 4(1):94–101, January 1959. CODEN LI-OCAH. ISSN 0024-3590.

Doty:1957:EPS

[DO57] Maxwell S. Doty and Mikihiko Oguri. Evidence for a photo synthetic daily periodicity. Limnology and Oceanography, 2(1):37–40, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1956:AEL

[EAP56] W. T. Edmondson, G. C. Anderson, and Donald R. Peterson. Artificial eutrophication of Lake Washington. Limnology and Oceanography, 1(1):47–53, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Ellis:1957:RRF

[EG57] Robert J. Ellis and Howard Gowing. Relationship retween food supply and condition of wild brown trout, Salmo trutta Linnaeus, in a Michigan stream. Limnology and Oceanography, 2(4):299–308, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Emery:1956:DSI

[Eme56] K. O. Emery. Deep standing internal waves in California Basins. Limnology and Oceanography, 1(1):35–41, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Emery:1958:BBS

[Eme58] K. O. Emery. Bacterial bottom sampler. Limnology and Oceanography, 4(2):109–111, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Eicher:1957:ELF

[ER57] George J. Eicher, Jr. and George A. Rounsefell. Effects of lake fertilization by volcanic activity on abundance of salmon. Limnology and Oceanography, 2(2):70–76, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Essenwanger:1957:HWM

[Ess57] Oskar Essenwanger. Handbuch der wellen der meere und ozeane. Limnology and Oceanography, 2(1):44, January 1957. CODEN LI-OCAH. ISSN 0024-3590.

Eyster:1958:BWC

[Eys58] Clyde Eyster. Bioassay of water from a concretion-forming marl lake. Limnology and Oceanography, 3(4):455–458, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Farris:1959:CEG

[Far59] David A. Farris. A change in the early growth rates of four larval marine fishes. Limnology and Oceanography, 4(1):29–36, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Fingerman:1958:EFI

[FLM58] Milton Fingerman, Mildred E. Lowe, and William C. Mobberly, Jr. Environmental factors involved in setting the phases of tidal rhythm of color change in the fiddler crabs Uca pugilator and Uca minax. Limnology and Oceanography, 3(3):271–282, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Fukai:1959:TAM

[FM59] Rinnosuke Fukai and W. W. Meinke. Trace analysis of marine organisms: a comparison of activation analysis and conventional methods. Limnology and Oceanography, 4(4):398–408, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Friedman:1956:DBL

[FNCR56] Irving Friedman, Daniel R. Norton, Douglas B. Carter, and Alfred C. Redfield. The deuterium balance of Lake Maracaibo. Limnology and Oceanography, 1(4):239–246, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Frey:1956:SPP

[Fre56] David G. Frey. Symposium on primary production: Preface. Limnology and Oceanography, 1(2):71, April 1956. CODEN LIOCAH. ISSN 0024-3590.

Frey:1957:AIC

[Fre57a] D. G. Frey. Aquatic insects of California. Limnology and Oceanography, 2(1):44–46, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Frey:1957:BMA

[Fre57b] D. G. Frey. E. A. Birge. A memoir; with an appraisal of Birge the limnologist: an explorer of lakes. Limnology and Oceanography, 2 (1):43, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Frey:1957:MAP

[Fre57c] D. G. Frey. A manual of aquatic plants. Limnology and Oceanography, 3(1):297, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Frey:1958:BRE

[Fre58a] D. G. Frey. Book review: Ergebnisse der deutschen limnologischen Venezuela-Expedition 1952. Bd. I. Limnology and Oceanography, 3 (2):242–243, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Frey:1958:KBS

[Fre58b] D. G. Frey. A key to the British species of freshwater Cladocera. Limnology and Oceanography, 3(4):482–483, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Geissbuhler:1957:HEG

[Gei57] Hans Geissbühler. Hydrobotanik. I. Energiehaushalt (German) [Hydrobotany. I. Energy budget]. Limnology and Oceanography, 2(1): 44, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1957:CCL

[Gor57a] Eville Gorham. The chemical composition of lake waters in Halifax County, Nova Scotia. Limnology and Oceanography, 2(1):12–21, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1957:CCS

[Gor57b] Eville Gorham. The chemical composition of some natural waters in the Cairn Gorm–Strath Spey District of Scotland. Limnology and Oceanography, 2(2):143–154, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1957:ICS

[Gor57c] Eville Gorham. The ionic composition of some lowland lake waters from Cheshire, England. Limnology and Oceanography, 2(1):22–27, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1958:OFB

[Gor58a] Eville Gorham. Observations on the formation and breakdown of the oxidized microzone at the mud surface in lakes. Limnology and Oceanography, 3(3):291–298, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1958:PLN

[Gor58b] Eville Gorham. The physical limnology of northern Britain: an epitome of the bathymetrical survey of the Scottish freshwater lochs, 1897–1909. Limnology and Oceanography, 4(2):40–50, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Gessner:1958:IOT

[GP58] Fritz Gessner and Federico Pannier. Influence of oxygen tension on respiration of phytoplankton. Limnology and Oceanography, 3(4): 478–480, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Guillard:1958:PEC

[GW58] Robert R. L. Guillard and Peter J. Wangersky. The production of extracellular carbohydrates by some marine flagellates. Limnology and Oceanography, 3(4):449–454, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1958:LWSa

[HA58] F. R. Hayes and E. H. Anthony. Lake water and sediment: I. Characteristics and water chemistry of some Canadian East Coast lakes. Limnology and Oceanography, 3(3):299–307, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1959:LWSb

[HA59] F. R. Hayes and E. H. Anteiony. Lake water and sediment: VI. The standing crop of bacteria in lake sediments and its place in the classification of lakes. Limnology and Oceanography, 4(3):299–315, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Hayne:1956:BPI

[HB56] Don W. Hayne and Robert C. Ball. Benthic productivity as influenced by fish predation*. Limnology and Oceanography, 1(3): 162–175, July 1956. CODEN LIOCAH. ISSN 0024-3590.

Hartman:1956:BC

[HE56] Olga Hartman and K. O. Emery. Bathypelagic coelenterates. Limnology and Oceanography, 1(4):304–312, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Hartman:1957:SCD

[HE57] Olga Hartman and K. O. Emery. Simplified computations for the dynamic height method of current determination in lakes. Limnology and Oceanography, 1(4):155–157, April 1957. CODEN LI-OCAH. ISSN 0024-3590.

Henson:1959:EIW

[Hen59] E. B. Henson. Evidence of internal wave activity in Cayuga Lake, New York. Limnology and Oceanography, 4(4):441–447, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Hepher:1958:DPA

[Hep58] B. Hepher. On the dynamics of phosphorus added to fishponds in Israel. Limnology and Oceanography, 4(2):84–100, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Hasler:1958:SOH

[HHWB58] A. D. Hasler, R. M. Horrall, W. J. Wisby, and W. Braemer. Sunorientation and homing in fishes. Limnology and Oceanography, 3 (4):353–361, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Hicks:1959:PON

[Hic59] Steacy D. Hicks. The physical oceanography of Narragansett Bay. Limnology and Oceanography, 4(3):316–327, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1959:LWSa

[HM59] F. R. Hayes and M. A. Macaulay. Lake water and sediment: V. Oxygen consumed in water over sediment cores. Limnology and Oceanography, 4(3):291–298, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1958:LWSc

[HP58] F. R. Hayes and J. E. Phillips. Lake water and sediment: IV. Radiophosphorus equilibrium with mud, plants, and bacteria under oxidized and reduced conditions. Limnology and Oceanography, 3 (4):459–475, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1958:LWSb

[HRC58] F. R. Hayes, B. L. Reid, and M. L. Cameron. Lake water and sediment: II. Oxidation-Reduction relations at the mud–water interface. Limnology and Oceanography, 3(3):308–317, July 1958. CO-DEN LIOCAH. ISSN 0024-3590.

Hulburt:1957:DNA

[Hul57] Edward M. Hulburt. The distribution of Neomysis americana in the estuary of the Delaware River. Limnology and Oceanography, 2 (1):1–11, January 1957. CODEN LIOCAH. ISSN 0024-3590.

John:1956:OSF

[JH56] Kenneth R. John and Arthur D. Hasler. Observations on some factors affecting the hatching of eggs and the survival of young shallow-water Cisco, Leucichthys artedi LeSueur, in Lake Mendota, Wisconsin. Limnology and Oceanography, 1(3):176–194, July 1956. CODEN LIOCAH. ISSN 0024-3590.

Jannasch:1959:BPS

[JJ59a] Holger W. Jannasch and Galen E. Jones. Bacterial populations in sea water as determined by different methods of enumeration. Limnology and Oceanography, 4(2):128–139, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Jones:1959:ABS

[JJ59b] Galen E. Jones and Holger W. Jannasch. Aggregates of bacteria in sea water as determined by treatment with surface active agents. Limnology and Oceanography, 4(3):269–276, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Kanwisher:1959:POE

[Kan59] John Kanwisher. Polarographic oxygen electrode. Limnology and Oceanography, 4(2):210–217, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Krumholz:1959:SKK

[KC59] Louis A. Krumholz and Gerald A. Cole. Studies on a Kentucky Knobs Lake. IV: Some limnological conditions during an unusually cold winter. Limnology and Oceanography, 4(4):367–385, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Ketchum:1958:PAS

[Ket58] B. H. Ketchum. Proceedings of the American Society of Limnology, and Oceanography, Inc. Limnology and Oceanography, 4(2):114– 117, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Ketchum:1959:PAS

[Ket59] Bostwick H. Ketchum. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 4(1): 115–117, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1957:POP

[KH57] Alan J. Kohn and Philip Helfrich. Primary organic productivity of a Hawaiian coral reef. Limnology and Oceanography, 3(1):241–251, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Kutkuiin:1958:NPN

[Kut58] Joseph H. Kutkuiin. Notes on the precision of numerical and volumetric plankton estimates from small-sample concentrates. Limnology and Oceanography, 4(2):69–83, January 1958. CODEN LI-OCAH. ISSN 0024-3590.

Livingstone:1958:EAE

[LBL58] D. A. Livingstone, Kirk Bryan, Jr., and R. G. Leahy. Effects of an Arctic environment on the origin and development of freshwater lakes. Limnology and Oceanography, 3(2):192–214, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Lewin:1958:MTA

[Lew58] Ralph A. Lewin. The mucilage tubes of Amphipleura rutilans. Limnology and Oceanography, 4(2):111–113, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Lund:1959:SCC

[Lun59] J. W. G. Lund. A simple counting chamber for nannoplankton. Limnology and Oceanography, 4(1):57–65, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Mackereth:1958:PCS

[Mac58] F. J. H. Mackereth. A portable core sampler for lake deposits. Limnology and Oceanography, 3(2):181–191, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Mann:1958:AFS

[Man58] K. H. Mann. Annual fluctuations in sulphate and bicarbonate hardness in ponds. Limnology and Oceanography, 3(4):418–422, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Mccombie:1959:SRB

[Mcc59] A. M. Mccombie. Some relations between air temperatures and the surface water temperatures of lakes. Limnology and Oceanography, 4(3):252–258, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Mcleod:1957:ECP

[Mcl57] G. C. Mcleod. The effect of circularly polarized light on the photosynthesis and chlorophyll-a synthesis of certain marine algae. Limnology and Oceanography, 2(4):360–362, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Mcneill:1959:AAA

[Mcn59] John Mcneill. Antibacterial activity of Antarctic marine phytoplankton. Limnology and Oceanography, 4(4):419–424, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Miller:1958:EWW

[Mil58] Arthur R. Miller. The effects of winds on water levels on the New England coast. Limnology and Oceanography, 3(1):1–14, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Mclachlan:1959:UCF

[MM59] Jack Mclachlan and Guy C. Macleod. The use of conversion factors for the determination of the concentration of nutrients in culture medial. Limnology and Oceanography, 4(2):218–219, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Montgomery:1957:OOP

[Mon57] R. B. Montgomery. Oceanic observations of the Pacific: 1949. Limnology and Oceanography, 1(4):160, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Montgomery:1958:EVL

[Mon58] R. B. Montgomery. Errata in vol. 2 of Limnology and Oceanography. Limnology and Oceanography, 2(2):118, January 1958. CODEN LI-OCAH. ISSN 0024-3590.

Montgomery:1959:TC

[Mon59] R. B. Montgomery. Townsend Cromwell, 1922–1958. Limnology and Oceanography, 4(2):228–229, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Mchugh:1959:LWA

[MOP59] J. L. Mchugh, R. T. Oglesby, and A. L. Pacheco. Length, weight, and age composition of the menhaden catch in Virginia waters. Limnology and Oceanography, 4(2):145–162, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Mortimer:1959:BRT

[Mor59] C. H. Mortimer. Book review: A treatise on limnology. Volume 1. Geography, physics and chemistry. Limnology and Oceanography, 4 (1):108–114, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Mcconnell:1959:CPM

[MS59] William J. Mcconnell and William F. Sigler. Chlorophyll and productivity in a mountain river. Limnology and Oceanography, 4(3): 335–351, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Macpherson:1958:LWS

[MSH58] L. B. Macpherson, N. R. Sinclair, and F. R. Hayes. Lake water and sediment: III. The effect of pH on the partition of inorganic phosphate between water and oxidized mud or its ash. Limnology and Oceanography, 3(3):318–326, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1959:EWC

[Mur59] Garth I. Murphy. Effect of water clarity on albacore catches. Limnology and Oceanography, 4(1):86–93, January 1959. CODEN LI-OCAH. ISSN 0024-3590.

Nicholls:1959:SAM

[NCB59] G. D. Nicholls, Herbert Curl, Jr., and V. T. Bowen. Spectrographic analyses of marine plankton. Limnology and Oceanography, 4(4): 472–478, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Odum:1956:PPF

[Odu56] Howard T. Odum. Primary production in flowing waters. Limnology and Oceanography, 1(2):102–117, April 1956. CODEN LIOCAH. ISSN 0024-3590.

Odum:1957:PPM

[Odu57] Howard T. Odum. Primary production measurements in eleven Florida springs and a marine turtle-grass community. Limnology and Oceanography, 2(2):85–97, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Ohle:1956:BPE

[Ohl56] Waldemar Ohle. Bioactivity, production, and energy utilization of lakes. Limnology and Oceanography, 1(3):139–149, July 1956. CODEN LIOCAH. ISSN 0024-3590.

Odum:1958:UPP

[OKB58] Eugene P. Odum, Edward J. Kuenzler, and Sister Marion Xavier Blunt. Uptake of P^{32} and primary productivity in marine benthic algae. Limnology and Oceanography, 3(3):340–348, July 1958. CO-DEN LIOCAH. ISSN 0024-3590.

Patterson:1958:TSS

[Pat58] Ashton M. Patterson. Turbulence spectrum studies in the sea with hot wires. Limnology and Oceanography, 3(2):171–180, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Pratt:1959:TSE

[PB59] David M. Pratt and Harold Berkson. Two sources of error in the oxygen light and dark bottle method. Limnology and Oceanography, 4(3):328–334, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Pratt:1956:EFA

[PC56] David M. Pratt and Donald A. Campbell. Environmental factors affecting growth in Venus mercenaria. Limnology and Oceanography, 1(1):2–17, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Penfound:1956:PPV

[Pen56] Wm. T. Penfound. Primary production of vascular aquatic plants. Limnology and Oceanography, 1(2):92–101, April 1956. CODEN LIOCAH. ISSN 0024-3590.

Pennak:1957:SCL

[Pen57] Robert W. Pennak. Species composition of limnetic zooplankton communities. Limnology and Oceanography, 3(1):222–232, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Pierce:1958:CIW

[Pie58] E. Lowe Pierce. The Chaetognatha of the inshore waters of North Carolina. Limnology and Oceanography, 3(2):166–170, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Posgay:1958:OSS

[PN58] J. A. Posgay and K. Duane Norman. An observation on the spawning of the sea scallop, Placopecten magellanicus (Gmelin), on Georges Bank. Limnology and Oceanography, 3(4):478, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1959:APS

[Pom59] Lawrence R. Pomeroy. Algal productivity in salt marshes of Georgia. Limnology and Oceanography, 4(4):386–397, April 1959. CO-DEN LIOCAH. ISSN 0024-3590.

Pratt:1959:PNB

[Pra59] David M. Pratt. The phytoplankton of Narragansett Bay. Limnology and Oceanography, 4(4):425–440, April 1959. CODEN LI-OCAH. ISSN 0024-3590.

Proctor:1957:SAA

[Pro57] Vernon W. Proctor. Studies of algal antibiosis using Haematococcus and Chlamydomonas. Limnology and Oceanography, 2(2):125–139, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Prey:1958:MPP

[PS58] D. G. Prey and J. B. Stahl. Measurements of primary production on Southampton Island in the Canadian Arctic. Limnology and Oceanography, 3(2):215–221, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Prowse:1958:SGS

[PT58] G. A. Prowse and J. F. Talltng. The seasonal growth and succession of plankton algae in the white Nile. Limnology and Oceanography, 3(2):222–238, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1956:ODB

[PTR56] Lawrence R. Pomeroy, Harold H. Tiaskin, and Robert A. Ragotzkie. Observations on dinoflagellate blooms. Limnology and Oceanography, 1(1):54–60, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Raymont:1958:SMC

[RA58] J. E. G. Raymont and M. N. E. Adams. Studies on the mass culture of Phaeodactylum. Limnology and Oceanography, 3(2):119–136, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Rawson:1956:AIT

[Raw56] D. S. Rawson. Algal indicators of trophic lake types. Limnology and Oceanography, 1(1):18–25, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Raymont:1959:RSP

[Ray59] J. E. G. Raymont. The respiration of some planktonic copepods: III. The oxygen requirements of some American species. Limnology and Oceanography, 4(4):479–491, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Richards:1956:SOA

[RC56] Francis A. Richards and Nathaniel Corwin. Some oceanographic applications of recent determinations of the solubility of oxygen in sea water. Limnology and Oceanography, 1(4):263–267, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Region:1958:ASP

[RD58] Lake Erie Region and Charles C. Davis. An approach to some problems of secondary production in the western. Limnology and Oceanography, 3(1):15–28, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Redfield:1956:A

[Red56] Alfred C. Redfield. Announcement. Limnology and Oceanography, 1(1):1, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Reid:1957:BHA

[Rei57] George K. Reid. Biologic and hydrographic adjustment in a disturbed Gulf Coast Estuary. Limnology and Oceanography, 3(1): 198–212, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Reid:1958:CDG

[Rei58] Joseph L. Reid, Jr. A comparison of drogue and GEK measurements in deep water. Limnology and Oceanography, 3(2):160–165, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Rice:1956:AES

[Ric56] T. R. Rice. The accumulation and exchange of strontium by marine planktonic algae. Limnology and Oceanography, 1(2):123–138, April 1956. CODEN LIOCAH. ISSN 0024-3590.

Ricker:1958:ECM

[Ric58a] W. E. Ricker. Evolution and classification of the mountain caddisflies. Limnology and Oceanography, 3(4):481, October 1958. CO-DEN LIOCAH. ISSN 0024-3590.

Ricker:1958:KAN

[Ric58b] W. E. Ricker. A key to the adults and nymphs of British stoneflies (Plecoptera). Limnology and Oceanography, 3(4):481, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Riley:1957:PNC

[Ril57] Gordon A. Riley. Phytoplankton of the North Central Sargasso Sea, 1950–52. Limnology and Oceanography, 3(1):252–270, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1959:LAM

[RM59] John H. Ryther and David W. Menzel. Light adaptation by marine phytoplankton. Limnology and Oceanography, 4(4):492–497, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Rittenberg:1958:CBS

[RMI58] Sydney C. Rittenberg, Tod Mittwer, and Daniel Iyler. Coliform bacteria in sediments around three marine sewage outfalls. Limnology and Oceanography, 4(2):101–108, January 1958. CODEN LIOCAH. ISSN 0024-3590.

Rounsefell:1959:DEFa

[Rou59a] George A. Rounsefell. On the dynamics of exploited fish populations. Limnology and Oceanography, 3(1):230–231, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Rounsefell:1959:DEFb

[Rou59b] George A. Rounsefell. On the dynamics of exploited fish populations. Limnology and Oceanography, 3(1):231–233, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Ragotzkie:1957:LHD

[RP57] Robert A. Ragotzkie and Lawrence R. Pomeroy. Life history of a dinoflagellate bloom. Limnology and Oceanography, 2(2):62–69, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Ravera:1956:BSN

[RT56] Oscar Ravera and Vittorio Tonolli. Body size and number of eggs in diaptomids, as related to water renewal in mountain lakes. Limnology and Oceanography, 1(2):118–122, April 1956. CODEN LI-OCAH. ISSN 0024-3590.

Rice:1959:UAL

[RW59] T. R. Rice and Virginia M. Willis. Uptake, accumulation and loss of radioactive cerium-144 by marine planktonic algae. Limnology and Oceanography, 4(3):277–290, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1957:EPP

[RY57] J. H. Ryther and C. S. Yentsch. The estimation of phytoplankton production in the ocean from chlorophyll and light data. Limnology and Oceanography, 3(1):281–286, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1958:PPC

[RY58] John H. Ryther and Charles S. Yentsch. Primary production of continental shelf waters off New York. Limnology and Oceanography, 3(3):327–335, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1956:MPP

[Ryt56a] John H. Ryther. The measurement of primary production. Limnology and Oceanography, 1(2):72–84, April 1956. CODEN LIOCAH. ISSN 0024-3590.

Ryther:1956:POF

[Ryt56b] John H. Ryther. Photosynthesis in the ocean as a function of light intensity. Limnology and Oceanography, 1(1):61–70, January 1956. CODEN LIOCAH. ISSN 0024-3590.

Saur:1956:HBB

[SA56] J. F. T. Saur and E. R. Anderson. The heat budget of a body of water of varying volume. Limnology and Oceanography, 1(4): 247–251, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Saila:1956:EMS

[Sai56] Saul B. Saila. Estimates of the minimum size-limit for maximum yield and production of chain pickerel, Esox niger LeSueur, in Rhode Island. Limnology and Oceanography, 1(3):195–201, July 1956. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1958:BSB

[San58] Howard L. Sanders. Benthic studies in Buzzards Bay. I. Animal– sediment relationships. Limnology and Oceanography, 3(3):245–258, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Schaefer:1958:PIO

[SB58] Milner B. Schaefer and Yvonne M. M. Bishop. Particulate iron in offshore waters of the Panama Bight and in the Gulf of Panama. Limnology and Oceanography, 3(2):137–149, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Swain:1959:SDA

[SBM59] F. M. Swain, A. Blumentals, and R. Millers. Stratigraphic distribution of amino acids in peats from Cedar Creek Bog, Minnesota, and Dismal Swamp, Virginia. Limnology and Oceanography, 4(2): 119–127, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Shapiro:1957:CBS

[Sha57] Joseph Shapiro. Chemical and biological studies on the yellow organic acids of lake water. Limnology and Oceanography, 2(3):161– 179, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Shimada:1958:DFP

[Shi58] Bell M. Shimada. Diurnal fluctuation in photosynthetic rate and chlorophyll “a” content of phytoplankton from Eastern Pacific Waters. Limnology and Oceanography, 3(3):336–339, July 1958. CO-DEN LIOCAH. ISSN 0024-3590.

Starr:1957:ECG

[SJ57] Theodore J. Starr and Mary Elizabeth Jones. The effect of copper on the growth of bacteria isolated from marine environments. Limnology and Oceanography, 2(1):33–36, January 1957. CODEN LIOCAH. ISSN 0024-3590.

Starr:1957:PVB

[SJM57] Theodore J. Starr, Mary E. Jones, and Domingo Martinez. The production of vitamin B_{12}-active substances by marine bacteria. Limnology and Oceanography, 2(2):114–119, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Smayda:1957:PSL

[Sma57] Theodore J. Smayda. Phytoplankton studies in Lower Narragansett Bay. Limnology and Oceanography, 2(4):342–359, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Starr:1957:VBA

[SMF57] Theodore J. Starr, Domingo Martinez, and Walter Fosberg. The vitamin-B_{12} activity of mullet and shark serum. Limnology and Oceanography, 2(2):111–113, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Slobodkin:1956:ERF

[SR56] L. Basil Slobodkin and Sumner Richman. The effect of removal of fixed percentages of the newborn on size and variability in populations of Daphnia pulicaria (Forbes). Limnology and Oceanography, 1(3):209–237, July 1956. CODEN LIOCAH. ISSN 0024-3590.

StJohn:1958:VSZ

[St.58] Philip A. St. John. A volumetric study of zooplankton distribution in the Cape Hatteras Area. Limnology and Oceanography, 3(4): 387–397, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Steen:1958:DSO

[Ste58] H. Steen. Determinations of the solubility of oxygen in pure water. Limnology and Oceanography, 3(4):423–426, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Sullivan:1958:PF

[Sul58] Charlotte M. Sullivan. The physiology of fishes. Limnology and Oceanography, 3(3):350–352, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Sutcliffe:1957:IMD

[Sut57] W. H. Sutcliffe, Jr. An improved method for the determination of preserved plankton volumes. Limnology and Oceanography, 3(1): 295–296, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Talling:1957:SOS

[Tal57] J. F. Talling. Some observations on the stratification of Lake Victoria. Limnology and Oceanography, 3(1):213–221, July 1957. CODEN LIOCAH. ISSN 0024-3590.

Turner:1957:TMD

[TB57] Harry J. Turner, Jr. and David L. Belding. The tidal migrations of Donax variabilis say. Limnology and Oceanography, 2(2):120–124, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Turekian:1957:SVS

[Tur57] Karl K. Turekian. The significance of variations in the strontium content of deep sea cores. Limnology and Oceanography, 2(4):309– 314, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1959:NWM

[Tyl59] John E. Tyler. Natural water as a monochromator. Limnology and Oceanography, 4(1):102–105, January 1959. CODEN LIOCAH. ISSN 0024-3590.

Vallentyne:1956:ECP

[Val56] J. R. Vallentyne. Epiphasic carotenoids in post-glacial lake sediments. Limnology and Oceanography, 1(4):252–262, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1956:PPL

[Ver56] Jacob Verduin. Primary production in lakes. Limnology and Oceanography, 1(2):85–91, April 1956. CODEN LIOCAH. ISSN 0024-3590.

Verduin:1957:DVP

[Ver57] Jacob Verduin. Daytime variations in phytoplankton photosynthesis. Limnology and Oceanography, 2(4):333–336, October 1957. CO-DEN LIOCAH. ISSN 0024-3590.

Weiss:1957:RSF

[WB57] Charles M. Weiss and James L. Botts. The response of some freshwater fish to isopropyl methylphosphonofluoridate (Sarin) in water. Limnology and Oceanography, 2(4):363–370, October 1957. CODEN LIOCAH. ISSN 0024-3590.

Whaley:1958:SSP

[Wha58] Richard C. Whaley. A submersible sampling pump1: Notes and comment. Limnology and Oceanography, 3(4):476–477, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1956:FIC

[Wie56] Wolfgang Wieser. Factors influencing the choice of substratum in Cumella vulgaris Hart (Crustacea, Cumacea). Limnology and Oceanography, 1(4):274–285, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1958:MS

[Wie58] W. Wieser. Meer und Strand. Limnology and Oceanography, 3(3): 349, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1959:EGS

[Wie59a] Wolfgang Wieser. The effect of grain size on the distribution of small invertebrates inhabiting the beaches of Puget Sound. Limnology and Oceanography, 4(2):181–194, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Wieser:1959:NSN

[Wie59b] Wolfgang Wieser. A note on subterranean nematodes from Chesapeake Bay, Md. Limnology and Oceanography, 4(2):225–227, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Williams:1958:SFS

[Wil58] Austin B. Williams. Substrates as a factor in shrimp distribution. Limnology and Oceanography, 3(3):283–290, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Wohlschlag:1959:SCB

[WJ59] Donald E. Wohlschlag and Rogelio O. Juliano. Seasonal changes in bluegill metabolism. Limnology and Oceanography, 4(2):195–209, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Wood:1958:PEA

[Woo58] R. D. Wood. Le phytoplancton, état actuel de nos connaissances sur les grands lacs est-africains et leur phytoplancton. (French) [Phytoplankton, current state of our knowledge of the great East African lakes and their phytoplankton]. Limnology and Oceanography, 3(3): 350, July 1958. CODEN LIOCAH. ISSN 0024-3590.

Whitney:1957:FCI

[WP57] Lester V. Whitney and Richard L. Pierce. Factors controlling the input of electrical energy into a fish (Cyprinus carpio L.) in an electrical field. Limnology and Oceanography, 2(2):55–61, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Wright:1958:LCF

[Wri58] John C. Wright. The limnology of Canyon Ferry Reservoir: I. Phytoplankton–zooplankton relationships in the euphotic zone during September and October, 1956. Limnology and Oceanography, 3 (2):150–159, April 1958. CODEN LIOCAH. ISSN 0024-3590.

Wright:1959:LCF

[Wri59] John C. Wright. Limnology of Canyon Ferry Reservoir: II. Phytoplankton standing crop and primary production. Limnology and Oceanography, 4(3):235–245, April 1959. CODEN LIOCAH. ISSN 0024-3590.

Whittaker:1957:OFS

[WV57] J. R. Whittaker and J. R. Vallentyne. On the occurrence of free sugars in lake sediment extracts. Limnology and Oceanography, 2 (2):98–110, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1956:SFA

[YD56] Charles S. Yentsch and Alyn C. Duxbury. Some of the factors affecting the calibration number of the Clarke-bumpus quantitative plankton sampler. Limnology and Oceanography, 1(4):268–273, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Yount:1956:FCS

[You56] James L. Yount. Factors that control species numbers in Silver Springs, Florida. Limnology and Oceanography, 1(4):286–295, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1957:STV

[YR57] C. S. Yentsch and J. H. Ryther. Short-term variations in phytoplankton chlorophyll and their significance. Limnology and Oceanography, 2(2):140–142, April 1957. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1958:PNO

[YV58] Charles S. Yentsch and Ralph F. Vaccaro. Phytoplankton nitrogen in the oceans. Limnology and Oceanography, 3(4):443–448, October 1958. CODEN LIOCAH. ISSN 0024-3590.

Zicker:1956:PRB

[ZBH56] E. L. Zicker, K. C. Berger, and A. D. Hasler. Phosphorus release from bog lake muds. Limnology and Oceanography, 1(4):296–304, October 1956. CODEN LIOCAH. ISSN 0024-3590.

Zobell:1959:TCA

[Zob59] Claude E. Zobell. Thermal changes accompanying the compression of aqueous solutions to deep-sea conditions. Limnology and Oceanography, 4(4):463–471, April 1959. CODEN LIOCAH. ISSN 0024-3590.